What is kates age if 3 times her age in 4 years time is 57
Answer by ankor@dixie-net.com(22740)

What is Kate's age if 3 times her age in 4 years time is 57
:
let a = her age now
write an equation for exactly what it says
3(a+4) = 57
divide both sides by 3
a + 4 = 57/3
a + 4 = 19
a = 19-4
a = 15 yr is her age now
:
;
Check it in the original equation